# Modeling phase transformations in Mn-rich disordered rocksalt cathodes with charge-informed machine-learning interatomic potentials

### v1

@@ III Results
To investigate the structural evolution, we performed charge-informed molecular dynamics from the disordered structure to monitor the phase change to the -phase. The initial (disordered) structure was generated by sampling from the equilibrium ensemble of CE-MC simulations at K. The MC structure represents the pristine (lithiated) composition (Li1.1Mn0.6Ti0.1O1.9F0.1) with a supercell containing 320 atoms. Based on the pristine structure, a delithiated structure (LiMn0.6Ti0.1O1.9F0.1, where represents cation vacancy) was then obtained by randomly removing Li ions and relaxing the delithiated configuration using the Product-CHGNet with a force convergence criterion of eV/Å. After relaxation, MD simulations were subsequently performed for ns with a time step of fs under the NVT ensemble at K. The MD trajectories were collected for subsequent structural analysis.
@@ III.1 Structure ordering analysis
Figure 4a presents the simulated X-ray diffraction (XRD) patterns derived from the structures within the time interval of to ns. To analyze the evolution of cation ordering on the FCC lattice, the atomic positions from the MD snapshots were refined to their nearest ideal Wyckoff sites. The gray dashed lines in Fig. 4a denote the characteristic peaks of spinel ordering (18∘, 35∘, etc.). At the onset of the MD simulation ( ns), there are no obvious peaks indicating spinel ordering. After approximately ns, the intensity of the spinel peaks increases and becomes more significant in subsequent time scales as the MD simulation proceeds, suggesting the formation of the spinel-like -phase.
Figure 4b illustrates the energy landscape across the simulated MD trajectories. The black points denote the potential energies, while the colored points represent the total energies. The energies are depicted relative to their values at the initial state. A notable decrease in energy (approximately 60 meV/anion) is observed for the -phase following a ns MD simulation. This energy reduction indicates that spinel-like cation ordering is thermodynamically favored at the composition Li0.6Mn0.8Ti0.1O1.9F0.1 when sufficient cation vacancies are present. This thermodynamic preference for spinel-like ordering aligns with previous computational studies that demonstrated the stability of spinel structures in Li-Mn-O systems, such as Li0.5MnO2 [30].
To study the change in local cation ordering, we analyzed the short-range order (SRO) indicated by the tetrahedron occupancy as a function of MD simulation time. Figure 4c presents the occurrence of 0-TM (solid green line) and 1-TM tetrahedral (solid purple line) channels in the MD structure, as a ratio to all interstitial tetrahedral sites. The 0-TM channel has a low migration barrier that kinetically favors fast Li-ion transport [31]. To compare with disordered phases, the dashed green lines in Fig. 4c indicate the random limit of the 0-TM channel for two Li contents ( for Li1.0 and for Li1.1 at the fully lithiated level) for reference. At the beginning of the MD simulation ( ns), the occurrence of 0-TM channels is approximately 0.03. This value is significantly lower than that of the random limit of Li1.1TMO2 (0.09), where the value is estimated based on a rocksalt structure with 0.1 excess Li and fully random cation ordering. The low value of 0-TM content in the initial structure indicates reduced percolating Li content and suggests that SRO is unfavorable for long-range Li percolation, which is consistent with previous predictions in Mn-Ti-based DRXs [28, 29]. As the phase transformation progresses, the occurrence of 0-TM channels increases, reaching approximately 0.06 after ns of MD simulation. This rise in the number of 0-TM channels agrees with the high 0-TM content in the spinel structure, indicating the emergence of spinel-like ordering.
@@ III.2 Charge distribution with ion rearrangement
To understand the correlation between Mn migration and its heterovalent state [9, 32], we inferred atomic charge states from the magnetic moments predicted by CHGNet. Figure 5a presents the site-wise distribution of magnetic moments for Mn ions. The heterovalent states of Mn are categorized as follows: Mn, Mn, and Mn. At the beginning of the MD simulation, Mn predominantly exists in valence states of +3 and +4. As the phase transformation progresses, the concentration of Mn2+ gradually increases, as shown by the increased frequency of the highlighted circles in Fig. 5a, particularly after the emergence of spinel-like ordering from 0.6 to 2.0 ns.
For a detailed description of Mn migration conditioned on its valence state, the occupancy of Mn and Mn per f.u. is plotted in Fig. 5b as an indication of hopping events to tetrahedral sites. Although previous nudged elastic band calculations reveal that Mn2+ has a lower migration barrier [11, 9], the atomic charges predicted by CHGNet show that most of the tetrahedral Mn are Mn from to ns during the MD simulation. This suggests that Mn3+ may migrate without the need for charge disproportionation () as proposed by Reed et al. [11]. The rise in Mn appears after around ns, which corresponds to the formation of spinel characteristic peaks (18∘, 35∘) in Fig. 4a. This evidence suggests that Mn3+ migration initiates the transformation before long-range spinel-like order is established. The subsequent emergence of tetrahedral Mn2+ appears to be a consequence of this long-range ordering (ns-scale) rather than a prerequisite for local cation hops (ps-scale). The simulation results are consistent with previous MD simulations of the LixMnO2 phase transformation revealed by Deng et al. [21] using the GGA-based MLIP.
@@ III.3 Intercalation voltage profile of -phase
To evaluate the effect of partial disorder on the electrochemical performance of the transformed structure from MD, we computed the intercalation voltage profiles using Product-CHGNet. We focused on two characteristic reactions in an ordered spinel LiMnO: (1) a 0-TM conversion reaction at a lower voltage, where the occupancies of Li ions are transformed from the octahedral sites to the tetrahedral sites collectively:
and (2) a conversion at a higher voltage, where Li is intercalated on the site. It is important to note that Eq. (3) is defined for an independent 0-TM channel (i.e., one that does not share Li with another tetrahedral channel). In an ordered LiMnO2 spinel, each Lioct is face-sharing with two 0-TM channels, where removing one Li from such a 0-TM channel results in a high defect energy [18]. The delithiation is, therefore, a collective conversion that exhibits a two-phase reaction between Li0.5MnO2 and LiMnO2 with a characteristic 3 V voltage plateau [33].
@@
Using Eq. (3), we enumerated different Li-vacancy configurations with fixed TMs and an anion framework (topotactically) at each (de)lithiation level based on the MD structure. The Product-CHGNet was used to perform structure relaxation for both lithiation and delithiation paths (see Appendix C for details). The energies of these (de)lithiated structures were used to construct the convex hull and the corresponding voltage profile as shown in Fig. 6a and Fig. 6b (see Methods).
@@
For the region with high Li concentration (), the 0-TM to conversion exhibits nearly solid-solution behavior, as evidenced by the convex hull being connected by several ground states in Fig. 6a. Particularly, for the early stage of delithiation (), the voltage exhibits a flat plateau with a minor slope, indicating that the delithiation through Eq. (3) occurs in similar local chemical environments and exhibits similar chemical potentials (voltages). We found that the voltage at is around 2.40 V, which is slightly lower than the experimental value. This discrepancy is likely attributable to the residual self-interaction error common in DFT functionals, even for r2SCAN, when a Hubbard U correction is not applied [34, 35]. Although the SCAN-based functional removes more self-interaction than previous LDA and GGA functionals [36], it does not fully remove it and underestimates the intercalation voltage, which is consistent with previous reports [18, 17].
@@
Upon reaching , lithium occupies all tetrahedral 0-TM channels (Li) in addition to the octahedral Li not in the 0-TM channels. Two competing delithiation processes are available, involving either the removal of Li or Li. We computed both pathways and found that the low-energy configurations near the convex hull within the range correspond to the removal of Li, while the high-energy configurations (represented by dark red markers in Fig. 6a) correspond to the removal of Li. These results indicate that even in a structure with partial spinel ordering, Li has a preference for delithiation before Li. In contrast, the delithiation in spinel LixMnO2 for entirely corresponds to the removal/insertion of Li with a V plateau. We also observe a voltage increase of V around (the end of 0-TM to Li conversion). The magnitude is not as large as that in spinel LiMnO2 ( V) or in the spinel-ordered LixMn0.8Ti0.1O1.9F0.1 (see Fig. 7c), suggesting the absence of two-phase reactions in the -phase and less inhomogeneity that could lead to battery performance degradation.
@@
To further demonstrate the effect of partial disorder, we evaluated the intercalation voltage profiles of Li1.1Mn0.8Ti0.1O1.9F0.1 in both the DRX and spinel phases for comparison. The DRX structure was selected from the structure derived from the CE-MC simulations (i.e., the configuration at s in the MD simulation). The spinel Li1.1Mn0.8Ti0.1O1.9F0.1 was generated through another CE-MC simulation without exchanges between the and sites, where all sites are occupied by Li1.0 and sites are occupied by Mn0.8, Ti0.1, and Li0.1. The MC simulation allows site exchange within sites and anion sites, which is designed to capture the underlying SRO of anions and cations in DRXs (e.g., Li-F SRO) [37]. The intercalation at high Li content in both DRX and spinel phases only included the 0-TM conversion reaction as described in Eq. (3).
@@
Figure 7a and 7b illustrate the formation energy convex hull and the voltage profile of the DRX phase, with the -phase plotted as a red line for comparison. Due to the insufficient number of 0-TM channels in the DRX structure, a relatively limited Li capacity from the 0-TM conversion is achieved (approximately 0.3 per f.u., as denoted by the green dashed line in Fig. 7a and 7b. The absence of spinel-like ordering results in a profile marked by a mildly steep increase as delithiation progresses. Notably, the overall voltage profile between of the DRX phase is higher than that of the -phase, suggesting a restricted amount of accessible capacity within an operational voltage window, particularly the capacity around V [9, 8]. In contrast, Figure 7b depicts the calculated results for spinel LixMn0.8Ti0.1O1.9F0.1. The voltage profile clearly displays a two-phase reaction between Li1.1 and Li0.6 with a linear convex hull in Fig. 7c. Due to the absence of disorder, the spinel phase exhibits a lower average voltage with a two-phase reaction between , with a voltage step more than 1 V presented at the end of 0-TM conversion. For , the intercalation corresponds to Li, which shows a similar high-voltage feature as the -phase. The effect of cation disorder on the spinel structure demonstrated in our simulations is in full agreement with the previous findings on the cation disorder effects in LiMnO2 spinel by Chen et al. [18].
@@ IV Discussion
In the evolving landscape of modern battery materials, the complexity of components and configurations exemplified by DRX cathode materials necessitates advanced computational models. Particularly, Mn-rich DRXs exhibit a phase transformation to a partially disordered spinel phase (-phase) via in situ electrochemical cycling [9, 8, 38] or chemical delithiation [39]. This transformation is challenging for conventional modeling approaches to capture, particularly the kinetic pathways of TM migration in such a complex oxide. While ab initio molecular dynamics could theoretically address this issue, its computational cost renders it impractical for probing large systems or long timescales. Alternatively, kinetic Monte Carlo (kMC) methods with CE models have been proposed to model cation ordering changes in both cathodes and solid electrolytes [40, 41]. However, parameterizing CEs for disordered oxides/oxyfluorides and accurately calculating migration barriers in diverse local environments remains a non-trivial endeavor [42].
@@
To address these challenges, we fine-tuned a charge-informed foundation MLIP with sampled r2SCAN-DFT calculated energies, forces, stresses, and magnetic moments. The multi-step PES sampling captures the complex interactions from diverse chemical environments and heterovalent states of TMs. Our MD simulations using the fine-tuned CHGNet successfully captured the phase transition from a disordered to a spinel-like configuration in Li0.6Mn0.8Ti0.1O1.9F0.1.
@@
The simulations reveal that the formation of the -phase promotes the occurrence of 0-TM channels, which play a crucial role in facilitating long-range Li-ion transport, whereas the DRX phase exhibits limited availability of these channels. This result agrees with the improved high-rate performance observed in experiments [9, 8, 39]. By analyzing the electrochemical intercalation voltage profiles of the -phase and comparing them to those of the DRX and spinel phases, two conclusions can be drawn from the simulation results: (1) The presence of partial cation disorder in the -phase effectively eliminates the two-phase reaction, as evidenced by the absence of a sharp voltage increase after Li0.5 removal (see Fig. 7c). (2) The introduction of spinel-like ordering increases the 0-TM occurrence (as indicated by the increased number of 0-TM channels in Fig. 4c) and thus expands the available Li capacity from the 0-TM to Li conversion reaction around 3 V (denoted by the green dashed line in Fig. 6b). The Li capacity from 0-TM to Li conversion is even higher than that of the spinel in Fig. 5c since the 0-TM channels in partially disordered phases are less correlated with each other compared to the spinel, where each Li is shared by two 0-TM channels.
@@
Although this study represents an integrated investigation of phase transformation in complex oxides combining MC simulations for configurational thermodynamics and MLIP-MD simulations for kinetic modeling, several limitations were identified that could be further addressed in future studies. For example, while our MLIP enables nanosecond-scale or million-atoms simulations with ab initio accuracy [43], many important physical processes—such as the formation of anti-phase boundaries in the -phase [39] or phase segregation in LixFePO4 cathodes [44]—occur on timescales still beyond this reach. Achieving ab initio accuracy while enabling simulations beyond typical MD timescales (such as using generative models for dynamic trajectories) holds promise for investigating these processes [45, 46, 47]. Regarding model fine-tuning, recent advances such as active learning with Bayesian force fields [48], stratified sampling from MLIP-driven MD [49], and diffusion-based generative models for mapping reaction pathways [50] provide more efficient sampling techniques beyond the passive sampling implemented in this work, which could enhance MLIP training accuracy by exploring the PES more effectively.
@@
In summary, we demonstrated the application of charge-informed MLIP to model the complex phase transformation from DRX to spinel-like -phase in Li-Mn-Ti-O-F systems. The MLIP-MD simulations revealed the mechanism of spinel-like ordering formation and its effect on electrochemical properties. As computational capabilities and new methodologies emerge–particularly in accelerating MLIP simulations and in generative modeling–these advances will pave the way for more accurate and efficient modeling of phase transformations in next-generation energy storage materials.

### v2

@@ III Results
To investigate the structural evolution as a disordered rocksalt transforms, we performed charge-informed molecular dynamics from the disordered structure to monitor the phase change to the -phase. The initial (disordered) structure was generated by sampling from the equilibrium ensemble of CE-MC simulations at K. The MC structure represents the pristine (lithiated) composition (Li1.1Mn0.8Ti0.1O1.9F0.1) with a supercell containing 320 atoms. Based on the pristine structure, a delithiated structure (LiMn0.8Ti0.1O1.9F0.1, where represents a cation vacancy) was then obtained by randomly removing Li ions and relaxing the delithiated configuration using the Product-CHGNet with a force convergence criterion of eV/Å. After relaxation, MD simulations were subsequently performed for ns with a time step of fs under the NVT ensemble at K. The MD trajectories were collected for subsequent structural analysis.
@@ III.1 Structure ordering analysis
Figure 4a presents the simulated X-ray diffraction (XRD) patterns derived from the structures within the time interval of to ns. To analyze the evolution of cation ordering on the FCC lattice, the atoms in the MD snapshots were moved to their nearest ideal Wyckoff sites. The gray dashed lines in Fig. 4a denote the characteristic peaks of spinel ordering (18∘, 35∘, etc.). At the onset of the MD simulation ( ns), there are no obvious peaks indicating spinel ordering. After approximately ns, the intensity of the spinel peaks increases and becomes more significant as the MD simulation proceeds, suggesting the formation of the spinel-like -phase.
Figure 4b illustrates the energy landscape across the simulated MD trajectories. The black points denote the potential energy, while the colored points represent the total energy. The energies are depicted relative to their values at the initial state. A notable decrease in energy (approximately 60 meV/anion) is observed after ns when the -phase has formed. This energy reduction indicates that spinel-like cation ordering is thermodynamically favored at the composition Li0.6Mn0.8Ti0.1O1.9F0.1 when sufficient cation vacancies are present. This thermodynamic preference for spinel-like ordering aligns with previous computational studies that demonstrated the stability of spinel structures in Li-Mn-O systems, such as Li0.5MnO2 [33].
To study the change in local cation ordering, we analyzed the short-range order (SRO) indicated by the tetrahedron occupancy as a function of MD simulation time. Figure 4c presents the occurrence of 0-TM (solid green line) tetrahedra in the MD structure, as a ratio to all tetrahedral sites. To compare with disordered phases, the dashed green lines in Fig. 4c indicate the random limit of the 0-TM tetrahedra for two Li contents ( for Li1.0 and for Li1.1 at the fully lithiated level) for reference. At the beginning of the MD simulation ( ns), the occurrence of 0-TM tetrahedra is approximately 0.03. This value is significantly lower than that of the random limit for Li1.1TMO2 (0.09). The low value of 0-TM content in the initial structure indicates that SRO is unfavorable for long-range Li percolation, which is consistent with previous predictions in Mn-Ti-based DRXs [29, 30]. As the phase transformation progresses, the occurrence of 0-TM tetrahedra increases, reaching approximately 0.06 after ns of MD simulation. This rise in the number of 0-TM tetrahedra agrees with the high 0-TM content in the spinel structure, indicating the emergence of spinel-like ordering.
@@ III.2 Charge distribution with ion rearrangement
To understand the correlation between Mn migration and its valence state [9, 34], we inferred atomic charge states from the magnetic moments predicted by CHGNet. Figure 5a presents the site-wise distribution of magnetic moments for Mn ions. The valence states of Mn are categorized as follows: Mn, Mn, and Mn. At the beginning of the MD simulation, Mn predominantly exists in valence states of +3 and +4, consistent with the partial delithiation of the material. As the phase transformation progresses, the concentration of Mn2+ gradually increases, as shown by the increased frequency of the highlighted circles in Fig. 5a, particularly after the emergence of spinel-like ordering from 0.6 to 2.0 ns.
For a detailed description of Mn migration conditioned on its valence state, the occupancy of Mn and Mn per f.u. is plotted in Fig. 5b. Although previous NEB calculations reveal that Mn2+ has the lower migration barrier of all Mn valence states [11, 9], the atomic charges predicted by CHGNet show that most of the tetrahedral Mn are Mn from to ns during the MD simulation. This suggests that Mn3+ may migrate without the need for charge disproportionation () as proposed by Reed et al. [11]. The rise in Mn appears after around ns, which corresponds to the formation of spinel characteristic peaks (18∘, 35∘) in Fig. 4a. This evidence suggests that Mn3+ migration initiates the transformation before long-range spinel-like order is established. The subsequent emergence of tetrahedral Mn2+ appears to be a consequence of this long-range ordering rather than a prerequisite for local cation hops. The simulation results are consistent with previous MD simulations of the LixMnO2 phase transformation revealed by Deng et al. [21] using the GGA-based MLIP.
@@ III.3 Intercalation voltage profile of -phase
To evaluate the effect of partial disorder on electrochemical performance, we computed the intercalation voltage profiles of the transformed -phase structure using Product-CHGNet. Specifically, we focus on the voltage characteristic of ordered spinel LixMn2O4 at lower voltage (for ), in which the occupancies of Li ions are transformed collectively from the tetrahedral sites to the empty octahedral sites upon lithiation. The reason this happens collectively in a well-ordered spinel is that (tet) and (oct) sites cannot be simultaneously occupied. Hence, upon lithiation the system empties the sites and occupies . However, in a (partially) disordered rocksalt structure, the , , and crystallographic labels cannot be defined precisely.
For this reason, we define a “0-TM” site as any interstitial tetrahedral site that is exclusively face-shared by four non-transition metal cations (i.e., Li or vacancy). This definition captures the essential local environment of the site in an ordered spinel which also has no transition metal neighbors. It also generalizes to disordered rocksalts where the connectivity of 0-TM sites is essential for evaluating the percolating Li content [10]. Using this more general description, we can expect two key intercalation processes:
(a) A lower-voltage reaction whereby a tetrahedral Li is removed and its octahedral sites become filled with Li. For such an isolated 0-TM site, this reaction can be written as:
(b) A higher-voltage reaction corresponding to the direct insertion of Li into an empty 0-TM tetrahedral site:
This framework allows us to consistently compare the electrochemical behavior of the ordered spinel, disordered rocksalt, and the intermediate -phase.
@@
We generated a series of Li-vacancy configurations to simulate the electrochemical (de)lithiation process topotactically. Starting with the fully lithiated structure obtained from MD, we held the TM and anion sublattices fixed. At each (de)lithiation level, distinct Li-vacancy orderings were created by selecting different 0-TM sites within the host structure and applying the local transformation described in Eq. (3) and (4). The Product-CHGNet was used to perform structure relaxation for both lithiation and delithiation paths (see Appendix C). The energies of these (de)lithiated structures were used to construct the convex hull and the corresponding voltage profile as shown in Fig. 6a and Fig. 6b.
@@
For the region with high Li concentration (), the 0-TM to conversion exhibits nearly solid-solution behavior, as evidenced by the convex hull being connected by several ground states in Fig. 6a. Particularly, for the early stage of delithiation (), the voltage exhibits a pseudo plateau with minor slope, indicating that the delithiation through Eq. (3) occurs in similar local chemical environments and exhibits similar chemical potentials (voltages). The computed voltage at is around 2.40 V, which is slightly lower than the experimental value. This discrepancy is likely attributable to the residual self-interaction error common in DFT functionals, even for r2SCAN, when a Hubbard U correction is not applied [35, 36]. Although the SCAN-based functional removes more self-interaction than previous LDA and GGA functionals [37], it does not fully remove it and underestimates the intercalation voltage, which is consistent with previous reports [18, 17].
@@
Upon reaching , lithium occupies all tetrahedral 0-TM sites (Li) in addition to the octahedral Li not in the 0-TM sites. Two competing delithiation processes are available, involving either the removal of Li or Li. We computed both pathways and found that the low-energy configurations near the convex hull in the range correspond to the removal of Li, while the high-energy configurations (represented by dark red markers in Fig. 6a) correspond to the removal of Li. These results indicate that even in a structure with partial spinel ordering, Li has a preference for delithiation before Li. In contrast, the delithiation in ordered spinel LixMnO2 for entirely corresponds to the removal/insertion of Li with a V plateau. We also observe a voltage increase of V around (the end of 0-TM to Li conversion). The magnitude is not as large as that in spinel LiMn2O4 ( V) or in the spinel-ordered LixMn0.8Ti0.1O1.9F0.1 (see Fig. 7d), suggesting the absence of two-phase reactions in the -phase.
@@
To further demonstrate the effect of partial disorder, we evaluated the intercalation voltage profiles of Li1.1Mn0.8Ti0.1O1.9F0.1 in both the DRX and spinel phases as a comparison. The DRX structure was selected from the CE-MC simulations (i.e., the configuration at s in the MD simulation). The spinel Li1.1Mn0.8Ti0.1O1.9F0.1 was generated through a separate CE-MC simulation without exchanges between the and sites, so that all sites are occupied by Li1.0 and sites are occupied by Mn0.8, Ti0.1, and Li0.1, similar to a Li-excess lithiated spinel. The MC simulation allows site exchange within sites and anion sites, which is designed to capture the underlying SRO of anions and cations in DRXs (e.g., Li-F SRO) [38]. The intercalation at high Li content in both DRX and spinel phases only included the 0-TM site conversion reaction as described in Eq. (3).
@@
Figure 7a and 7b illustrate the formation energy convex hull and the voltage profile of the DRX phase, with the -phase plotted as a red line for comparison. As there are few 0-TM sites in the DRX structure, a relatively limited Li capacity from the 0-TM conversion is achieved (approximately 0.3 per f.u., as denoted by the vertical green dashed line in Fig. 7a and 7b. The absence of spinel-like ordering results in a profile marked by a mildly steep voltage increase as delithiation progresses. Notably, the overall voltage profile of the DRX phase between is higher than that of the -phase, suggesting a restricted amount of accessible capacity around V [9, 8]. In contrast, Figure 7d depicts the calculated results for spinel LixMn0.8Ti0.1O1.9F0.1. The voltage profile clearly displays a two-phase reaction between Li1.1 and Li0.6 consistent with the linear convex hull in Fig. 7c. Due to the absence of disorder, the spinel phase exhibits a lower average voltage with a two-phase reaction between , and a voltage step larger than 1 V at the end of 0-TM conversion. For , the intercalation corresponds to Li, which shows a similar high-voltage feature as the -phase. The effect of cation disorder on the spinel structure demonstrated in our simulations is in agreement with the previous findings on the cation disorder effects in LiMnO2 spinel by Chen et al. [18].
@@ IV Discussion
In the evolving landscape of modern battery materials, the multirole of components and complexity of configurations exemplified by DRX cathode materials necessitates advanced computational models. Particularly, Mn-rich DRXs exhibit a phase transformation to a partially disordered spinel phase (-phase) via in situ electrochemical cycling [9, 8, 39] or chemical delithiation [40, 41]. This transformation is challenging for conventional modeling approaches, as it requires capturing the kinetic pathway of TM migration in the disordered environment. While ab initio molecular dynamics could theoretically address this issue, its computational cost renders it impractical for probing large systems or long timescales. Alternatively, kinetic Monte Carlo (kMC) methods with CE models have been proposed to model cation ordering changes in both cathodes and solid electrolytes [42, 43]. However, parameterizing CEs for disordered oxides/oxyfluorides and accurately calculating migration barriers in diverse local environments with DFT remains a non-trivial endeavor [44].
@@
To address these challenges, we fine-tuned a foundation MLIP with sampled r2SCAN-DFT calculated energies, forces, stresses, and magnetic moments. The multi-step PES sampling captures the complex interactions from diverse chemical environments and different valence states of TMs. Our MD simulations using the fine-tuned CHGNet successfully captured the phase transition from a disordered to a spinel-like configuration in Li0.6Mn0.8Ti0.1O1.9F0.1.
@@
The simulations reveal that the formation of the -phase from DRX promotes the increase of 0-TM configurations on tetrahedra, which play a crucial role in facilitating long-range Li-ion transport. This result is consistent with the improved high-rate performance observed in experiments [9, 8, 40]. By analyzing the simulated electrochemical intercalation voltage profiles of the -phase and comparing them to those of the DRX and spinel phases, two conclusions can be drawn: (1) The presence of partial cation disorder in the -phase effectively eliminates the two-phase reaction, as evidenced by the absence of a sharp voltage increase after Li0.5 removal (see Fig. 7d). This is consistent with earlier predictions on partially disordered spinels [18]. (2) The introduction of spinel-like ordering increases the 0-TM occurrence (as indicated by the increased number of 0-TM tetrahedra in Fig. 4c) and thus expands the available Li capacity from the 0-TM to Li conversion reaction around 3 V (denoted by the green dashed line in Fig. 6b). The Li capacity from 0-TM to Li conversion in the -phase is even higher than that of the spinel in Fig. 7d since the 0-TM sites in partially disordered phases are less spatially correlated with each other compared to the spinel, where each Li is shared by two 0-TM sites.
@@
Although this study represents an integrated investigation of phase transformation in complex oxides combining MC simulations for configurational thermodynamics and MLIP-MD simulations for kinetic modeling, several limitations were identified that could be further addressed in future studies. For example, while our MLIP enables nanosecond-scale or million-atoms simulations with ab initio accuracy [45], many important physical processes—such as the formation of anti-phase boundaries in the -phase [40] or phase segregation in LixFePO4 cathodes [46]—occur on timescales still beyond this reach. Achieving ab initio accuracy while enabling simulations beyond typical MD timescales (such as using generative models for dynamic trajectories) holds promise for investigating these processes [47, 48, 49]. Regarding model fine-tuning, recent advances such as active learning with Bayesian force fields [50], stratified sampling from MLIP-driven MD [51], and diffusion-based generative models for mapping reaction pathways [52, 53] provide more efficient sampling techniques beyond the passive sampling implemented in this work, which could enhance MLIP training accuracy by exploring the PES more effectively. Finally, it is important to clarify that the term ”charge” refers to the valence state (nominal charge) of the TM, which is inferred from the predicted magnetic moments. The current model does not explicitly compute long-range electrostatic (Coulombic) interactions. Looking forward, the model’s capabilities could be significantly enhanced by incorporating response charge prediction, such as Latent Ewald Summation (LES) [54]. LES integrates long-range electrostatics and the prediction of Born effective charges [55], which has demonstrated a universal augmentation for mainstream MLIPs [56]. We believe an LES-augmented CHGNet would yield a more physically robust description of charge, thereby improving the modeling of complex phenomena such as cation migration during phase transformations in oxides.
@@
In summary, we demonstrated the application of a MLIP to model the complex phase transformation from DRX to spinel-like -phase in Li-Mn-Ti-O-F systems. The MLIP-MD simulations revealed the mechanism of spinel-like ordering formation and its effect on electrochemical properties. As computational capabilities and new methodologies emerge–particularly in accelerating MLIP simulations and in generative modeling–these advances will pave the way for more accurate and efficient modeling of phase transformations in new energy storage materials.